OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5.

OpenStudy (souvik):

so..y=k/(2x) where k is some constant...put y=10 and x=3 and find k first

OpenStudy (anonymous):

y=k/2x for y=10, x=3. So 10=k/6 =>k=60 Equation becomes: y=60/2x=30/x now just put the values of x=5 to find corresponding y

OpenStudy (anonymous):

I still don't understand